# Non-Iterative Disentangled Unitary Coupled-Cluster based on Lie-algebraic structure

### v2

@@ IV.1 NI-DUCC-VQE versus ADAPT-VQE
Although the UsCCSDTQ-VQE method provides a systematic way to extend the UCC ansatz, and achieves desirable chemical accuracy for strongly correlated systems (see discussion in the previous Section), implementing triple and quadruple excitations for this method on quantum computers is costly, especially in terms of one- and two-qubit gates [25, 31, 30]. Therefore, in this Section, we will compare UsCCSDTQ-VQE with NI-DUCC-VQE in term of computational costs, including the number of CNOT gates, parameters, and function evaluations (also called optimization steps). Given that ADAPT-VQE algorithms have demonstrated significant improvements in term of accuracy, as well as a reduction in computational costs (see Table S1 in the supplementary materials [55]), we will focus on comparing the computational costs of NI-DUCC-VQE with fermionic-ADAPT-VQE [32], qubit-ADAPT-VQE [33], and QEB-ADAPT-VQE [35]. To ensure a fair comparison, we use data for LiH, H6, and BeH2 at bond distances of 1.546 Å, 1.5 Å, and 1.316 Å, respectively (as shown in Figure 6 of reference [35]). The numerical results were extracted from the link provided in [62]. Based on this data, we perform NI-DUCC calculations on the same molecules bond distances, and their estimate computational resources (see Figure 4).
@@
In Figure 4(a), at the $10^{-12}$ Hartree level of accuracy, where larger circuits are expected, the QEB-VQE-ADAPT method systematically outperforms both fermionic- and qubit-ADAPT-VQE in terms of CNOT efficiency. This superiority stems from the type of pool excitations used in the ansatz. According to [35], QEB-ADAPT-VQE employs efficient qubit excitation circuits, that are more effective than the fermionic excitations used in fermionic ADAPT-VQE. Moreover, qubit-ADAPT-VQE fails to surpass QEB-ADAPT-VQE energies at high levels of accuracy, simply because qubit evolutions enable local circuit optimizations, while the more rudimentary Pauli string evolutions used by qubit-ADAPT-VQE do not. Therefore it would require more iterations to approach the exact energy. The NI-DUCC CNOT counts exhibits similat results to QEB-ADAPT. This is due to the fact that the NI-DUCC ansatz elements consist in evolutions of $XY$-Pauli strings of length $p$ with an odd $Y$ Pauli terms, leading to CNOT gates scaling as $22p-$ 2 per Pauli string [33]. This indicates that the NI-DUCC excitations ansatz achieves an efficient scaling in the usage of CNOT gates.
@@
In terms of parameter counts, while NI-DUCC-VQE and QEB-ADAPT-VQE require significantly fewer CNOTs compared to fermionic-ADAPT-VQE, the QEB-ADAPT-VQE requires up to twice as many variational parameters, and the NI-DUCC-VQE requires up to three times as many (see Fig. 4(b)). This difference arises because NI-DUCC-VQE and QEB-ADAPT-VQE, assign one parameter to each qubit excitation (one Pauli string) and qubit excitation (single and double) evolutions in their ansatz, respectively, whereas fermionic-ADAPT-VQE assigns one parameter to a pair of spin-complement fermionic excitation evolutions. These spin-complement pairs can enforce parity conservation, which is not achieved by qubit excitation evolutions. Additionally, NI-DUCC requires more parameters due to the dynamic increase in the number of layers $k$ of the ansatz (see Eq. 3), even though the pool generators are complete and originate from the symmetric starter excitations, which conserves the parity symmetry property in NI-DUCC ansatz. The reader can refer to Table S1 in supplementary material [55] which compares the scales of NI-DUCC, UsCCSDTQ, and ADAPT-VQE versions in terms of CNOT and parameter counts.
@@
We also computed the number of function evaluations required by each of the studied algorithms, as shown in Figures 4(c) for H6 and 4(d) for BeH2. As noted in the computational procedure Section, all algorithms used the BFGS optimizer, with analytically calculated energy gradient vectors. In both benchmarks, we observe that the NI-DUCC-VQE algorithm requires only 800 function evaluations for H6, and 1684 for BeH2, to achieve an accuracy below 10-12 Hartree. In contrast, the ADAPT-VQE algorithms need at least 25 ×$\times 10^{3}$ to 30 ×$\times 10^{3}$ function evaluations to reach the same level of accuracy. This highlights the advantage of NI-DUCC-VQE in achieving full convergence, with a relatively small number of optimization steps, i.e. function evaluations. This efficiency is attributed to the combination of strongly symmetric correlated orbitals in the initial guess, with the closure algebraic relations, which accelerates the computations, and avoids false local minima, as observed in both molecules (See figures. 2 and 2 in Section III.1).
IV.2 Challenges in MCP generation as the number of qubits increases
@@
The main cost in constructing the NI-DUCC ansatz, lies in the precomputation step that involves building the symmetry-preserving MCP. The MCP consists of a set of generators that scale as $\mathcal{O}(n)$ with $n$ qubits, and can form a Lie algebraic closure set, as previously explained in the methodology. The definitions, properties, completeness conditions of MCP along with its symmetry-related conditions, are described in section 4 of the supplementary materials [55] (Comprehensive information about MCP generation can be also found in the reference [37]). The main steps involved in constructing the MCP and their associated computational costs on a classical computer, are as follows:
@@
The Python code for generating the MCP can be found in [60]. However, as mentioned in the computational details (See Section II), we transformed the Python code into a C++ program. Indeed, comparing the performance of Python and C++ implementations, there are several critical factors come into play in the context of High-Performance Computing (HPC): execution speed, memory consumption, and ease of optimization (see further explanations of these factors in section 5 of the supplementary materials [55]). In our tests, storing for example the FullGroup using python as a python set, is memory inefficient when ones tries to scale up the number of qubits. To overcome this issue, we designed a specific C++ data structure set, that minimizes memory consumption to the detriment of searching operations, in order to be able to hold larger sets and to increase the number of qubits. This grants the ability to run 14 qubits systems encompassing 7 electrons, on a local laptop, without suffering from memory issues. In practice, the C++ approach only requires 95 MiB instead of the 14 GiB required by the python code. Our C++ subroutine, is fully optimized, which enables to generate MCPs for up to 20 qubits while insuring sufficient large amount of memory. However, with the present implementation, the time required to perform product groups computations (i.e between 20 and beyond) remains very high. Table 1 provides an example of the cost associated with the lists generation for the FullSet, the FullGroup, and the full symmetric starters, for constructing the MCP. It is evident that these costs scale exponentially with $n$, in both computer memory and computational time (see Figures S3 (a) and S3 (b) of the supplementary materials [55]):
@@
The MCP generation corresponds actually to the final step of the NI-DUCC-VQE algorithm (see Section 2.I.2). Overall, the other steps are not particularly costly, as we only generate the starters of double excitations using sparsity conditions, and then impose symmetry preservation, on the qubit excitations of these starters. These steps, along with a symmetry-preserving MCP generation, rely solely on classical computation. Therefore, the NI-DUCC ansatz is dynamically constructed with multiple layers (see Eq. 3), without any “pre-circuit” measurements from the quantum computer. Indeed, once the increased load on a classical computer for MCP generation (beyond 20 qubits) is addressed, the cost of NI-DUCC-VQE would be very low. This is because our DUCC ansatz is fixed and non-iterative, requiring only the optimization of the ansatz parameters, which still necessitates measurements similar to the standard VQE process. In contrast, methods like ADAPT-VQE require substantial pre-circuit measurements, to obtain gradient measurements as well as iterative optimization parameters, for adaptively constructing their ansatz.
@@ V Conclusion
We have introduced the NI-DUCC-VQE algorithm, designed to construct a disentangled unitary coupled-cluster ansatz, based on a specific set of qubit excitation generators. This method scales linearly with the number of qubits, and leverages strong symmetry orbitals along with Lie algebraic properties. On a test set of several molecules, we have demonstrated that our method systematically improves accuracy, achieving full convergence, and reaching the Full Configuration Interaction (FCI) reference energy. This is particularly significant at larger internuclear distances, where increased electron correlation leads to multiple electron configurations contributing comparably to the wavefunction. We have highlighted that this aspect is frequently ignored in the current state-of-the-art quantum computing fixed ansatz for quantum chemistry. This is true even when using higher-order fermionic excitations in the unitary coupled cluster (UCC) method, such as up to triple excitations (UCCSDT) or even up to quadruple excitations (UsCCSDTQ). Such methods fail to capture all the correlations, especially when molecular bond lengths deviate from equilibrium. NI-DUCC not only demonstrates an high level of accuracy but also proves to be a hardware-efficient approach, making it well-suited for implementation on NISQ devices. The CNOT count for the NI-DUCC ansatz scales as $\mathcal{O}(knp)$ with its depth $k$, due to the $22n-$ 2 generators of MCP that scale linearly, which results in a relatively shallow circuit depth. When comparing NI-DUCC-VQE to ADAPT-VQE, particularly the promising QEB-ADAPT-VQE algorithm, which targets a shallower circuit depth, we found that similar CNOT counts can be achieved. NI-DUCC-VQE has also shown that with a small number of layers, it struggles to accurately find the exact ground state. However, with a sufficient number of layers, the algorithm consistently, and rapidly, converges to the FCI solution. We also observed that when the number of layers is sufficiently large, the energy decreases exponentially with each optimization step. This demonstrates that if the NI-DUCC ansatz incorporates more layers, the optimization hypersurface becomes more favorable. This effectively eliminates the false local minima, as demonstrated with the tested molecules in the article, thereby achieving FCI solution with fewer iterations. This finding paves the way for future discussions on using the NI-DUCC method to address the barren plateaus problem, where a cost function minimization reveals flat areas in the cost landscape of a variational quantum algorithm.
@@
Despite these key advantages, we have shown that the NI-DUCC approach comes with a notable cost due to the time complexity involved in the generation of the product group, which is a necessary step to validate the completeness condition of the pool operators. This limitation presently affects large-scale computations beyond 20 qubits. Addressing this issue will require further work. First, we are currently introducing the NI-DUCC approach in the Hyperion-1 GPU-accelerated high performance emulator [68] to benefit from further code optimization that will go further than the presently discussed C++ implementation. Second, it will be crucial to find better approaches related to the Lie algebraic properties for constructing optimal set of excitations. Nevertheless, despite our current limitations with increasing qubits, NI-DUCC-VQE holds potential for applications on NISQ devices. It can be used for active space selections to target large molecules with frozen core electrons, potentially reaching FCI solutions. Additionally, it could be applied to target excited states in both open and closed-shell molecules.

### v3

@@ IV.2 NI-DUCC-VQE versus ADAPT-VQE
Given that ADAPT-VQE algorithms have demonstrated significant improvements in term of accuracy, as well as a reduction in computational costs (see Table S1 in the Supplementary Materials [59]), in this subsection we will focus on comparing the computational costs of NI-DUCC-VQE with fermionic-ADAPT-VQE [32], qubit-ADAPT-VQE [33], and QEB-ADAPT-VQE [35]. To ensure a fair comparison, we use data for LiH, H6, and BeH2 at bond distances of 1.546 Å, 1.5 Å, and 1.316 Å, respectively (as shown in Figure 6 of reference [35]). The numerical results were extracted from the link provided in [67]. Based on this data, we perform NI-DUCC calculations on the same molecules bond distances, and their estimate computational resources (see Figure 4).
@@
In Figure 4(a), at the $10^{-12}$ Hartree level of accuracy, where larger circuits are expected, the QEB-VQE-ADAPT method systematically outperforms both fermionic- and qubit-ADAPT-VQE in terms of CNOT efficiency. This superiority stems from the type of pool excitations used in the ansatz. According to [35], QEB-ADAPT-VQE employs efficient qubit excitation circuits, that are more effective than the fermionic excitations used in fermionic ADAPT-VQE. Moreover, qubit-ADAPT-VQE fails to surpass QEB-ADAPT-VQE energies at high levels of accuracy, simply because qubit evolutions enable local circuit optimizations, while the more rudimentary Pauli string evolutions used by qubit-ADAPT-VQE do not. Therefore it would require more iterations to approach the exact energy. The NI-DUCC CNOT counts exhibit similar results to QEB-ADAPT. This is due to the fact that the NI-DUCC ansatz elements consist in evolutions of $XY$-Pauli strings of length $p$ with an odd $Y$ Pauli terms, leading to CNOT gates scaling as $22p-$ 2 per Pauli string [33]. This indicates that the NI-DUCC excitations ansatz achieves an efficient scaling in the usage of CNOT gates.
@@
In terms of parameter counts, while NI-DUCC-VQE and QEB-ADAPT-VQE require significantly fewer CNOTs compared to fermionic-ADAPT-VQE, the QEB-ADAPT-VQE requires up to twice as many variational parameters, and the NI-DUCC-VQE requires up to three times as many (see Fig. 4(b)). This difference arises because NI-DUCC-VQE and QEB-ADAPT-VQE, assign one parameter to each qubit excitation (one Pauli string) and qubit excitation (single and double) evolutions in their ansatz, respectively, whereas fermionic-ADAPT-VQE assigns one parameter to a pair of spin-complement fermionic excitation evolutions. These spin-complement pairs can enforce parity conservation, which is not achieved by qubit excitation evolutions. Additionally, NI-DUCC requires more parameters due to the dynamic increase in the number of layers $k$ of the ansatz (see Eq. 3), even though the pool generators are complete and originate from the symmetric starter excitations, which conserves the parity symmetry property in NI-DUCC ansatz. The reader can refer to Table S1 in Supplementary Materials [59] which compares the scales of NI-DUCC, UsCCSDTQ, and ADAPT-VQE versions in terms of CNOT and parameter counts.
@@
We also computed the number of function evaluations required by each of the studied algorithms, as shown in Figures 4(c) for H6 and 4(d) for BeH2. As noted in the computational procedure Section, all algorithms used the BFGS optimizer, with analytically calculated energy gradient vectors. In both benchmarks, we observe that the NI-DUCC-VQE algorithm requires only 800 function evaluations for H6, and 1684 for BeH2, to achieve an accuracy below 10-12 Hartree. In contrast, the ADAPT-VQE algorithms need at least 25 ×$\times 10^{3}$ to 30 ×$\times 10^{3}$ function evaluations to reach the same level of accuracy. This highlights the advantage of NI-DUCC-VQE in achieving full convergence, with a relatively small number of optimization steps, i.e. function evaluations. This efficiency is attributed to the combination of strongly symmetric correlated orbitals in the initial guess, with the closure algebraic relations, which accelerates the computations, and avoids false local minima, immune to barren plateaus, as observed in both molecules (see Figures 2 and 2 in Section III.1).
Overall, when considering algorithmic costs, it is crucial to assess the CNOT gate counts, the parameter counts and function evaluations. While fewer CNOT gates reduce the circuit depth and enhance robustness against noise, a higher parameter count can improve convergence rates, as observed in our simulations. Specifically, NI-DUCC-VQE has demonstrated rapid convergence, and achieving FCI solutions, with significantly fewer function evaluations compared to ADAPT-VQE methods. Therefore we can say while NI-DUCC may require more parameters, its efficient CNOT usage and rapid convergence demonstrate that this trade-off can lead to superior performance in practical quantum computing hardware implementation scenarios.
IV.3 Challenges in MCP generation as the number of qubits increases
@@
The main cost in constructing the NI-DUCC ansatz, lies in the precomputation step that involves building the symmetry-preserving MCP. The MCP consists of a set of generators that scale as $\mathcal{O}(n)$ with $n$ qubits, and can form a Lie algebraic closure set, as previously explained in the methodology. The definitions, properties, completeness conditions of MCP along with its symmetry-related conditions, are described in section 4 of the Supplementary Materials [59] (note that comprehensive information about MCP generation can be also found in the reference [37]). The main steps involved in constructing the MCP and their associated computational costs on a classical computer, are as follows:
@@
The time and memory required to generate and process the product group and Lie algebra through string multiplications grow exponentially with the number of qubits. We also refer readers to step d in Figure 1 for visual clarification and to Section 4 of the supplementary material [59] for a more detailed explanation.
In terms of coding, the Python code for generating the MCP can be found in reference [65]. However, as mentioned in the computational details (See Section II), we transformed the Python code into a C++ program. Indeed, comparing the performance of Python and C++ implementations, there are several critical factors come into play in the context of High-Performance Computing (HPC): execution speed, memory consumption, and ease of optimization (see further explanations of these factors in section 5 of the Supplementary Materials [59]). In our tests, storing for example the FullGroup using python as a python set, is memory inefficient when ones tries to scale up the number of qubits. To overcome this issue, we designed a specific C++ data structure set, that minimizes memory consumption to the detriment of searching operations, in order to be able to hold larger sets and to increase the number of qubits. This grants the ability to run 14 qubits systems encompassing 7 electrons, on a local laptop, without suffering from memory issues. In practice, the C++ approach only requires 95 MiB instead of the 14 GiB required by the python code. Our C++ subroutine, is fully optimized, which enables to generate MCPs for up to 20 qubits while insuring sufficient large amount of memory. However, with the present implementation, the time required to perform product groups computations (i.e between 20 and beyond) remains very high. Table 4 provides an example of the cost associated with the lists generation for the FullSet, the FullGroup, and the full symmetric starters, for constructing the MCP. It is evident that these costs scale exponentially with $n$, in both computer memory and computational time (see Figures S3 (a) and S3 (b) of the Supplementary Materials [59]):
@@
The MCP generation corresponds actually to the final step of the NI-DUCC-VQE algorithm (see Section 2.I.2). Overall, the other steps are not particularly costly, as we only generate the starters of double excitations using sparsity conditions, and then impose symmetry preservation, on the qubit excitations of these starters. These steps, along with a symmetry-preserving MCP generation, rely solely on classical computation. Therefore, the NI-DUCC ansatz is dynamically constructed with multiple layers (see Eq. 3), without any “pre-circuit” measurements from the quantum computer. Indeed, once the computational load for MCP generation beyond 20 qubits is managed, NI-DUCC’s fixed, non-iterative ansatz will keep costs low, requiring only standard VQE-style measurements. Unlike ADAPT-VQE, which incurs high costs from iterative gradient measurements, NI-DUCC leverages symmetry constraints with Lie algebraic conditions, which enables faster convergence and efficient scaling. Currently, the time cost for DI-DUCC at 18 qubits reflects our specific simulation setup and appears costly. However, direct comparisons across algorithms remain challenging due to differences in the followings: resource requirements, parameter scaling, and convergence behavior. As we work on scaling DI-DUCC simulations beyond 18 qubits, we expect that NI-DUCC’s unique advantages will allow it to outperform gradient-based methods like ADAPT-VQE in time cost, facilitating a comprehensive cross-algorithm comparison in future work.
@@ V Conclusion
We have introduced the NI-DUCC-VQE algorithm, designed to construct a disentangled unitary coupled-cluster ansatz, based on a specific set of qubit excitation generators. This method scales linearly with the number of qubits, and leverages strong symmetry orbitals along with Lie algebraic properties. On a test set of several molecules, we have demonstrated that our method systematically improves accuracy, achieving full convergence, and reaching the Full Configuration Interaction (FCI) reference energy. This is particularly significant at larger internuclear distances, where increased electron correlation leads to multiple electron configurations contributing comparably to the wavefunction. We emphasize that current state-of-the-art fixed ansatz methods in quantum computing for quantum chemistry, such as unitary coupled cluster (UCC) approaches with up to triple (UCCSDT) and quadruple (UsCCSDTQ) fermionic excitations, often fail to capture all correlations, especially as molecular bond lengths move away from equilibrium. In contrast, NI-DUCC not only achieves high accuracy but is also hardware-efficient, making it particularly suitable for NISQ devices. The CNOT count for the NI-DUCC ansatz scales as $\mathcal{O}(knp)$ with its depth $k$, due to the $22n-$ 2 generators of MCP that scale linearly, which results in a relatively shallow circuit depth. Our comparison of CNOT counts between NI-DUCC-VQE and advanced fixed ansatz methods, like COMPASS and COMPACT, reveals that NI-DUCC reduces CNOT usage by more than a factor of 3 for a given accuracy across tested molecules. Remarkably, NI-DUCC surpasses the accuracy of COMPASS and COMPACT with fewer CNOTs, reaching FCI solutions. Comparing NI-DUCC-VQE with ADAPT-VQE, especially QEB-ADAPT-VQE optimized for shallow circuits, we found similar CNOT counts to reach FCI solutions. Unlike ADAPT-VQE, NI-DUCC like COMPACT ansatz, it requires no precircuit measurements for operator gradients. Simulations show that with sufficient layering, NI-DUCC-VQE achieves exponential energy reduction per step, efficiently converging to the FCI solution by avoiding false local minima. This result could suggest NI-DUCC’s potential for addressing barren plateaus in variational algorithms. However, as NI-DUCC layers increase, its parameter count exceeds that of UsCCDTQ, QEB-ADAPT-VQE, COMPACT, and COMPASS, which highights the need for parameter compression strategies like those tools used in COMPASS for instance. This presents a promising direction for further future optimization in NI-DUCC.
@@
Despite these key advantages, we have shown that the NI-DUCC approach comes with a notable cost due to the time complexity involved in the generation of the product group, which is a necessary step to validate the completeness condition of the pool operators. This limitation presently affects large-scale computations beyond 20 qubits. Addressing this issue will require further work. First, we are currently introducing the NI-DUCC approach in the Hyperion-1 GPU-accelerated high performance emulator [72] to benefit from further code optimization that will go further than the presently discussed C++ implementation. Second, we are developing a machine learning model to accelerate the verification steps (outlined in Section IV.C (step 4)) by predicting whether a given set of starters meets the inseparability and completeness conditions(explained in details in section 4 of the supplementary Material [59]), thereby identifying optimal MCP candidates faster. Nevertheless, despite our current limitations with increasing qubits, NI-DUCC-VQE holds potential for applications on NISQ devices. It can be used for active space selections to target large molecules with frozen core electrons, potentially reaching FCI solutions. Using active space selections within 20 qubits, it can perform accurate calculations with larger chemical basis sets beyond the minimal STO-3G, effectively expanding its applicability to more complex quantum chemistry simulations as qubit resources grow. We aim to tackle the challenges associated with MCP generation as the number of qubits increases, thereby allowing NI-DUCC to advance further with larger basis sets. Additionally, it could be applied to target excited states in both open and closed-shell molecules.